exact value of f(1)
Problem 2. If f(x) = I+arctan () and f(0) = 1, what is the exact value of f(1)?
Answer #1

xfl t(X) Xtane tanca) sehel df (x) xtane tantu) xfl dr 2 ane tan(x) dx th + df (2) c27 x² tl Now, x dr dp 2 x+1=P =)2xdx=deane tan (3) x²ti 2 te 2 - 1x)) + = f(x) = en 182+11+ + Carctan (2) [C=Citez flode 2) flora en LP + 2 (tanta fo) = ln lotil t2. f(1) =ln2 + + 2 x R? 16 tl f(t) = In 2+ H² f (Ans) 32